OpenStudy (anonymous):

A certain wheel has a rotational inertia of 12 kg.m^2. As it turns through 5.0 rev its angular velocity increases from 5.0 rad/s to 6.0 rad/s. If the net torque is constant its value is ?

OpenStudy (anonymous):

when torque constant we used that eq : \[\omega=\tau (\Delta \Theta)\] \[\tau = (6/10\pi)\] so, Torque =1.89 but the correct answer was 2.1 N.m !!

OpenStudy (theeric):

This link might help: http://hyperphysics.phy-astr.gsu.edu/hbase/mi.html#mi

OpenStudy (theeric):

That equation can't be correct because it can be rearranged to be\[\tau =\frac{\omega [rad/s]}{\Delta\theta [rad]}\]That would mean torque is measured in \(1/s\) - also called \(Hz\) or Hertz.

OpenStudy (theeric):

So \(\tau = I\alpha\) could work for you (it's like \(F=ma\)), and \(\alpha\) is angular acceleration. Is there maybe a \(v_f^2=v_i^2+2ad\) sort of equation for rotation?

OpenStudy (theeric):

According to http://bama.ua.edu/~jharrell/PH105-S03/exercises/rot_mot_eqs.htm there is! \[\omega_f^2=\omega_i^2+2\alpha\left(\Delta \theta\right)\] And \(\Large\Delta\theta=5[rev]\times \frac{2\pi [rad]}{1[rev]}=10\pi [rad]\) I did the math, and I got \(2.1008452488130184321492656765172\) on the Windows calculator.

OpenStudy (anonymous):

Thank you very much :)

OpenStudy (theeric):

You're welcome! \(\Large\text{:)}\)